# Incentivizing Reasoning for Advanced Instruction-Following of Large Language Models

### v1

@@ Abstract
Existing large language models (LLMs) face challenges of following complex instructions, especially when multiple constraints are present and organized in paralleling, chaining, and branching structures. One intuitive solution, namely chain-of-thought (CoT), is expected to universally improve capabilities of LLMs. However, we find that the vanilla CoT exerts a negative impact on performance due to its superficial reasoning pattern of simply paraphrasing the instructions. It fails to peel back the compositions of constraints for identifying their relationship across hierarchies of types and dimensions. To this end, we propose a systematic method to boost LLMs in dealing with complex instructions via incentivizing reasoning for test-time compute scaling. First, we stem from the decomposition of complex instructions under existing taxonomies and propose a reproducible data acquisition method. Second, we exploit reinforcement learning (RL) with verifiable rule-centric reward signals to cultivate reasoning specifically for instruction following. We address the shallow, non-essential nature of reasoning under complex instructions via sample-wise contrast for superior CoT enforcement. We also exploit behavior cloning of experts to facilitate steady distribution shift from fast-thinking LLMs to skillful reasoners. Extensive evaluations on seven comprehensive benchmarks confirm the validity of the proposed method, where a 1.5B LLM achieves 11.74% gains with performance comparable to a 8B LLM. Codes and data are available at https://github.com/yuleiqin/RAIF.
@@ 1 Introduction
Large language models (LLMs) exhibited remarkable performance on real-world tasks  touvron2023llama ; grattafiori2024llama ; kojima2022large ; zhao2023survey . Such generalizability is built upon the instruction-following capabilities of LLMs ouyang2022training ; mu2023can ; zhou2023instruction . To benchmark whether LLMs can produce the desired outputs under complex instructions, existing studies predominantly focus on modeling various types of constraints and rules where all requirements are expected to be satisfied simultaneously zhang2024cfbench ; jing2023followeval ; he2024can ; zhou2023instruction . Recently, the compositions of constraints (And, Chain, Selection, and Nested) have been systemized wen2024benchmarking to further enhance the complexity of instructions and demonstrate that LLMs still fail to meet expectations under intricate structures.
To improve LLMs in solving complex instructions, most prior methods leverage two kinds of techniques: 1) supervised fine-tuning (SFT) xu2023wizardlm ; he2024complex , and 2) template-guided inference sun2024conifer ; chae2024language . Despite their effectiveness, these methods are all task-specific and prone to overfitting of limited types of instructions. For the former, a large amount of curated instruction-response pairs are required to ensure diversity. For the latter, it is almost impossible to exhaustively enumerate the templates for problem decomposition, reflection, and refinement beforehand. Therefore, it calls upon a scalable solution that is both constructive and generalizable without meticulous manual efforts.
One intuitive method is to directly apply chain-of-thought (CoT) reasoning wei2022chain on complex instructions, where LLMs might benefit from the free-form thinking for structure analysis with highlighted, valid constraints (see Fig. 1). However, our preliminary experiments show that such a prospective solution brings minimal or even negative performance gains, which contrasts strikingly against its effectiveness on maths problems. We observe that such a discrepancy arises from the underlying problem of "superficial" reasoning, where LLMs simply summarize the instructions briefly without developing the solid thinking upon the instruction itself. During such shallow, parroting-style reasoning, critical constraints and rules can be ignored and thereafter such misalignment leads to degraded performance. For maths problems, it is indispensable for LLMs to formalize step-by-step, divide-and-conquer process to achieve the final answer. On the contrary, for complex instructions, there exists no such nature or tendency of LLMs to forge deep reasoning as they are aligned to directly deliver responses without intermediate steps. Under such circumstance, we target at incentivizing reasoning that truly empowers LLMs with strategic planning and adherence to rules.
@@
In this paper, we propose to universally improve the instruction-following capabilities of LLMs under complex instructions. Inspired by the success of o1 jaech2024openai , R1 guo2025deepseek , and QwQ qwq32b , we resort to CoT for tackling complex instructions. We provide a practical guide for cultivation of effective reasoning tailored via reinforcement learning (RL). Specifically, our method addresses the following two main challenges: 1) the data scarcity of diverse complex instructions and 2) the secret recipe behind the formulation of effective CoT. For the former, although various data synthesis methods xu2023wizardlm ; wang2022self ; chung2024scaling ; kopf2023openassistant ; tunstall2023zephyr ; chen2024self ; luo2023wizardcoder ; sun2024conifer have been proposed to pile up instructions, they do not take into consideration the taxonomy of constraint types and compositions. In contrast, we stem from the atomic rules, constraints, and their compositions to perform self-evolving across various tasks and domains. Besides, our instruction scaling is integrated with our RL-based reasoning stimulation. Both rule-based and model-based verification approaches, which serve as reward sources later, are paired with the instructions during generation. For the latter, existing approaches that seek to reproduce R1-like reasoning hu2025openreasonerzeroopensourceapproach ; tinyzero ; deepscaler2025 ; zeng2025simplerl are all restricted to solving maths problems with pure rule-based rewards, which are inapplicable for our settings. We consequently keep an eye on the differences between maths and complex instructions, and propose our RL recipe with rule-centric reward modeling, enforcement on superior CoT, and control on policy distribution drift. Our contributions are three-fold:
@@ 2.1 Evaluation of Instruction Following of LLMs
Tremendous benchmarks have sprouted to specifically evaluate instruction following capabilities of LLMs in terms of semantic li2023alpacaeval ; zheng2023judging ; liu2023alignbench and format constraints he2024complex ; xia2024fofo ; jing2023followeval . Following the representative IFEval zhou2023instruction , recent studies extend the evaluation settings towards wild chat zhao2024wildchat ; lior2025wildifeval ; an2025ultraif , long context wu2024lifbench , multi-lingual li2024cif ; li2025xifbench ; dussolle2025m , multi-turn li2025structflowbench ; he2024multi , and multi-modal qian2024mia ; epstein2024mmmt scenarios. The hierarchical categorization haradacurse ; zhang2025iheval ; yang2024enhancing ; ziems2024tower and down-streaming applications yan2025codeif ; dong2024toward ; oh2024instructir are considered to collect fine-grained instruction for comprehensive evaluation.
In this paper, we measure the instruction following capabilities on the most common, comprehensive benchmarks including IFEval zhou2023instruction , ComplexBench wen2024benchmarking , CELLO he2024can , CFBench zhang2024cfbench , FB-Bench li2024fb , FollowBench jiang2023followbench , and InfoBench qin2024infobench . Different from studies that merely focus on the evaluation itself, we systematically investigate a general, scalable pipeline that enhances the instruction following consistently across models and benchmarks via test-time compute scaling.
@@ 2.2 Complex Rule and Constraint Following
Numerous studies enhanced rule and constraint following capabilities via data engineering qinunleashing . The top concerns are around the synthesis of complicated constraints without costly human intervention. Specifically, WizardLM xu2023wizardlm presents both in-breadth and in-depth evolving to generate complex instructions for tuning. Air liu2025air stems from the simple instruction distilled from documents and then iteratively add one or more constraints to increase complexity. Back-translation qi2024constraint is another popular technique to introduce additional constraints. Self-play dong2024self targets at generating constraints that can be verified by codes. SPAR cheng2024spar combines both self-play and tree-search for iterative refinement. The detailed taxonomy of rules and constraints zhang2024cfbench ; wang2024instructions ; sun2024conifer , together with their structures and relationship hayati2024chain ; wen2024benchmarking , allows precise synthesis control over instruction types. With respect to the post-training technique, most methods adopt the SFT sun2024beyond and RL huang2025musc ; wang2024direct ; zhang2024iopo ; wu2024meta ; ren2025step for preference alignment. Moreover, training-free workflows have also been proposed to decompose the instructions  ferraz2024llm and exploit verifiable feedback wang2025verifiable for refinement until all the check-boxes are ticked cook2024ticking .
@@ 2.3 Chain-of-Thought Reasoning
Recent progress on the chain-of-thought (CoT) wei2022chain ; zhang2024chain ; xu2025chain ; besta2024graph ; yao2023tree ; yang2024buffer has attracted attentions for advancing cognitive capabilities LLMs. Most of these approaches prompted LLMs to break down difficult questions into multiple small units to tackle them systematically before jumping to final answers kojima2022large ; press2022measuring ; arora2022ask ; zhou2022least ; khot2022decomposed ; zhang2022automatic . The emergence of OpenAI o1 jaech2024openai and DeepSeek R1 guo2025deepseek has incubated test-time compute scaling techniques for long CoT, where deep reasoning with extensive exploration and feasible reflection are encouraged chen2025towards . Both Monte Carlo Tree Search and RL with process and outcome rewards qin2024o1 ; huang2024o1 ; huang2025o1 are also emphasized in previous academic attempts to replicate slow-thinking LLMs. Compared with existing studies that focus on the mathematic and logic problem-solving, we are aimed at advancing instruction following capabilities. Motivated by R1 guo2025deepseek , we incentivize reasoning with group relative policy optimization (GRPO) shao2024deepseekmath and demonstrate that RL with rule-centric rewards ultimately pushes the limits of LLMs in following instructions.
@@ 3.2 Self-Evolving Instructions with Complex Rules and Constraints
We start by selecting a set of seed instructions $D_{\text{seed}}$ from the commonly used WildChat zhao2024wildchat and Alpaca alpaca datasets. To ensure the diversity of $D_{\text{seed}}$, we follow lu2023instag to tag each instruction by its topics and tasks for a wide-range selection of task abilities. Details on the tagging and selection process can be found in Sec. A.3.1.
@@
We adapt self-instruct wang2022self under different fine-grained rules and constraints zhang2024cfbench . In view of the verification techniques, both code-execution qiao2023making and LLM-as-a-Judge zheng2023judging are involved to provide evaluation feedback. For the former, we prepare a collection of constraint templates and their executable codes. Then, we randomly instantiate a combination of rules and constraints from the pool. To ensure their mutual compatibility, a pre-defined validity check is enforced to eliminate conflicts (e.g., The first paragraph must start with... and Wrap your entire reply with...). For the latter, we construct pairs of sub-instructions and scoring questions that stress on the style and semantic constraints. Such sub-instructions correspond to constraints with LLM-based evaluation (e.g., Is the answer written in the tone of Confucious?), which are complementary to those relying on code-based evaluation. An off-the-shelf LLM is utilized to perform few-shot in-breadth evolution. With respect to the composition of these atomic sub-instructions, we refer to he2024complex for definitions of And, Chain, and Selection. These sub-instructions are assembled for the integrated instructions. Details can be found in the Sec. A.3.2.
@@ 3.3 Reinforcement Learning for Reasoning Under Complex Instructions
We propose to incentivize reasoning of LLMs via RL ouyang2022training ; guo2025deepseek . The development of CoT is optimized towards being structured and sophisticated, which ultimately leads to improved final answers. Without loss of generality we adopt the GRPO shao2024deepseekmath algorithm. Given a query $x$ sampled from the distribution $D$ of complex instructions, the policy model $\pi_{\theta_{\text{old}}}$ from the previous iteration generates a group of $G$ individual outputs $\{y^{i}\}_{i=1}^{G}$. GRPO updates the policy $\pi_{\theta}$ by maximizing the objective:
@@
Compared with the proximal policy optimization (PPO) schulman2017proximal , the advantage of the $i$-th output response is computed by normalizing the group-level rewards $\{R^{i}\}_{i=1}^{G}$. PPO optimizes an additional critic model as value function via the generalized advantage estimation (GAE) DBLP:journals/corr/SchulmanMLJA15 . In consideration of the simplicity and our available computing resources, we employ the GRPO as our scale-up RL settings.
@@
To explicitly distinguish the reasoning from the answer contents, we employ a simple minimalist rule-based format reward that checks the existence of "<think>", "</think>", "<answer>" and "</answer>" tags. The format reward encourages the thinking contents encolosed solely between "<think>" and "</think>". It is noted that for R1-series reasoning models, "<answer>" and "</answer>" tags are not necessary for answer extraction.
$R^{i}_{\text{format}}=\left\{\begin{array}[]{lcl}+1&&{\textrm{if}\ \mathbbm{1}(\texttt{<think>}...\texttt{</think>}\texttt{<answer>}...\texttt{</answer>}\in y^{i}),}\\ -1&&{\textrm{otherwise.}}\end{array}\right.$ (4)
@@
$R^{i}_{\text{accuracy}}=\left\{\begin{array}[]{lcl}+2&&{\textrm{if}\ \text{is\_followed}(y^{i}|c_{j})\ \forall c_{j}\in x_{C}^{\text{active}},}\\ \sum_{j=1}^{C}\frac{\mathbbm{1}(c_{j}\in x_{C}^{\text{active}}\ \&\&\ \text{is\_followed}(y^{i}|c_{j}))}{\mathbbm{1}(c_{j}\in x_{C}^{\text{active}})}&&{\text{elif}\ \text{is\_followed}(y^{i}|c_{j})\ \exists c_{j}\in x_{C}^{\text{active}},}\\ -2&&{\textrm{otherwise.}}\end{array}\right.$ (5)
It is noted that the detailed implementation of the verification depends on the constraints zhang2024cfbench . For the lexical-level constraints (e.g., keywords and phrases), the numerical constraints (e.g., letters and words), and the format constraint (e.g., JSON, XML, LaTeX, HTML, and Markdown), we resort to simple heuristics with python which provides precise feedback zhou2023instruction ; dong2024self . In contrast, for the semantic-level constraints (e.g., themes and perspectives), stylistic constraint (e.g., writing styles, tones, and role-plays), linguistic constraints (e.g., dialects and morphologies), we bring in the reward model $r_{\phi}$ for judgment he2024complex ; qin2024infobench ; wen2024benchmarking . However, different from the original GRPO where $r_{\phi}$ implicitly scores the responses in terms of instruction following, we explicitly employ $r_{\phi}$ to check the following conditions of constraints. Besides, $r_{\phi}$ delivers scalar scoring in previous studies while we request for efficient boolean validation (True or False). Since multiple constraints might be active for steering generation, we develop the piecewise reward that promotes more constraint satisfaction while penalizing greatly the extreme cases. In total, our rule-centric reward is defined below.
@@
$\mathbbm{1}(\max(\{R^{i}_{\text{accuracy}}\}_{i=1}^{G})\geq\min(\{\hat{R}^{i}_{\text{accuracy}}\}^{G}_{i=1})\cdot\mathcal{J}_{\text{GRPO}}^{i}.$

### v2

@@ Abstract
Existing large language models (LLMs) face challenges of following complex instructions, especially when multiple constraints are present and organized in paralleling, chaining, and branching structures. One intuitive solution, namely chain-of-thought (CoT), is expected to universally improve capabilities of LLMs. However, we find that the vanilla CoT exerts a negative impact on performance due to its superficial reasoning pattern of simply paraphrasing the instructions. It fails to peel back the compositions of constraints for identifying their relationship across hierarchies of types and dimensions. To this end, we propose a systematic method to boost LLMs in dealing with complex instructions via incentivizing reasoning for test-time compute scaling. First, we stem from the decomposition of complex instructions under existing taxonomies and propose a reproducible data acquisition method. Second, we exploit reinforcement learning (RL) with verifiable rule-centric reward signals to cultivate reasoning specifically for instruction following. We address the shallow, non-essential nature of reasoning under complex instructions via sample-wise contrast for superior CoT enforcement. We also exploit behavior cloning of experts to facilitate steady distribution shift from fast-thinking LLMs to skillful reasoners. Extensive evaluations on seven comprehensive benchmarks confirm the validity of the proposed method, where a 1.5B LLM achieves 11.74% gains with performance comparable to a 8B LLM.
@@ 1 Introduction
Large language models (LLMs) exhibited remarkable performance on real-world tasks  Touvron et al. (2023); Grattafiori et al. (2024); Kojima et al. (2022); Zhao et al. (2023). Such generalizability is built upon the instruction-following capabilities of LLMs Ouyang et al. (2022); Mu et al. (2023); Zhou et al. (2023). To benchmark whether LLMs can produce the desired outputs under complex instructions, existing studies predominantly focus on modeling various types of constraints and rules where all requirements are expected to be satisfied simultaneously Zhang et al. (2024a); Jing et al. (2023); He et al. (2024a); Zhou et al. (2023). Recently, the compositions of constraints (And, Chain, Selection, and Nested) have been systemized Wen et al. (2024) to further enhance the complexity of instructions and demonstrate that LLMs still fail to meet expectations under intricate structures.
To improve LLMs in solving complex instructions, most prior methods leverage two kinds of techniques: 1) supervised fine-tuning (SFT) Xu et al. (2023); He et al. (2024b), and 2) template-guided inference Sun et al. (2024a); Chae et al. (2024). Despite their effectiveness, these methods are all task-specific and prone to overfitting of limited types of instructions. For the former, a large amount of curated instruction-response pairs are required to ensure diversity. For the latter, it is almost impossible to exhaustively enumerate the templates for problem decomposition, reflection, and refinement beforehand. Therefore, it calls upon a scalable solution that is both constructive and generalizable without meticulous manual efforts.
One intuitive method is to directly apply chain-of-thought (CoT) reasoning Wei et al. (2022) on complex instructions, where LLMs might benefit from the free-form thinking for structure analysis with highlighted, valid constraints (see Fig. 1). However, our preliminary experiments show that such a prospective solution brings minimal or even negative performance gains, which contrasts strikingly against its effectiveness on maths problems. We observe that such a discrepancy arises from the underlying problem of ”superficial” reasoning, where LLMs simply summarize the instructions briefly without developing the solid thinking upon the instruction itself. During such shallow, parroting-style reasoning, critical constraints and rules can be ignored and thereafter such misalignment leads to degraded performance. For maths problems, it is indispensable for LLMs to formalize step-by-step, divide-and-conquer process to achieve the final answer. On the contrary, for complex instructions, there exists no such nature or tendency of LLMs to forge deep reasoning as they are aligned to directly deliver responses without intermediate steps. Under such circumstance, we target at incentivizing reasoning that truly empowers LLMs with strategic planning and adherence to rules.
@@
In this paper, we propose to universally improve the instruction-following capabilities of LLMs under complex instructions. Inspired by the success of o1 Jaech et al. (2024), R1 Guo et al. (2025), and QwQ Qwen (2025), we resort to CoT for tackling complex instructions. We provide a practical guide for cultivation of effective reasoning tailored via reinforcement learning (RL). Specifically, our method addresses the following two main challenges: 1) the data scarcity of diverse complex instructions and 2) the secret recipe behind the formulation of effective CoT. For the former, although various data synthesis methods Xu et al. (2023); Wang et al. (2022a); Chung et al. (2024); Köpf et al. (2023); Tunstall et al. (2023); Chen et al. (2024); Luo et al. (2023); Sun et al. (2024a) have been proposed to pile up instructions, they do not take into consideration the taxonomy of constraint types and compositions. In contrast, we stem from the atomic rules, constraints, and their compositions to perform self-evolving across various tasks and domains. Besides, our instruction scaling is integrated with our RL-based reasoning stimulation. Both rule-based and model-based verification approaches, which serve as reward sources later, are paired with the instructions during generation. For the latter, existing approaches that seek to reproduce R1-like reasoning Hu et al. (2025); Pan et al. (2025); Luo et al. (2025); Zeng et al. (2025) are all restricted to solving maths problems with pure rule-based rewards, which are inapplicable for our settings. We consequently keep an eye on the differences between maths and complex instructions, and propose our RL recipe with rule-centric reward modeling, enforcement on superior CoT, and control on policy distribution drift. Our contributions are three-fold:
@@ 2.1 Evaluation of Instruction Following of LLMs
Tremendous benchmarks have sprouted to specifically evaluate instruction following capabilities of LLMs in terms of semantic Li et al. (2023); Zheng et al. (2023); Liu et al. (2023) and format constraints He et al. (2024b); Xia et al. (2024); Jing et al. (2023). Following the representative IFEval Zhou et al. (2023), recent studies extend the evaluation settings towards wild chat Zhao et al. (2024); Lior et al. (2025); An et al. (2025), long context Wu et al. (2024a), multi-lingual Li et al. (2024a, 2025a); Dussolle et al. (2025), multi-turn Li et al. (2025b); He et al. (2024c), and multi-modal Qian et al. (2024); Epstein et al. (2024) scenarios. The hierarchical categorization Harada et al. ; Zhang et al. (2025); Yang et al. (2024a); Ziems et al. (2024) and down-streaming applications Yan et al. (2025); Dong et al. (2024a); Oh et al. (2024) are considered to collect fine-grained instruction for comprehensive evaluation.
In this paper, we measure the instruction following capabilities on the most common, comprehensive benchmarks including IFEval Zhou et al. (2023), ComplexBench Wen et al. (2024), CELLO He et al. (2024a), CFBench Zhang et al. (2024a), FB-Bench Li et al. (2024b), FollowBench Jiang et al. (2023a), and InfoBench Qin et al. (2024a). Different from studies that merely focus on the evaluation itself, we systematically investigate a general, scalable pipeline that enhances the instruction following consistently across models and benchmarks via test-time compute scaling.
@@ 2.2 Complex Rule and Constraint Following
Numerous studies enhanced rule and constraint following capabilities via data engineering Qin et al. (2024b). The top concerns are around the synthesis of complicated constraints without costly human intervention. Specifically, WizardLM Xu et al. (2023) presents both in-breadth and in-depth evolving to generate complex instructions for tuning. Air Liu et al. (2025) stems from the simple instruction distilled from documents and then iteratively add one or more constraints to increase complexity. Back-translation Qi et al. (2024) is another popular technique to introduce additional constraints. Self-play Dong et al. (2024b) targets at generating constraints that can be verified by codes. SPAR Cheng et al. (2024) combines both self-play and tree-search for iterative refinement. The detailed taxonomy of rules and constraints Zhang et al. (2024a); Wang et al. (2024a); Sun et al. (2024a), together with their structures and relationship Hayati et al. (2024); Wen et al. (2024), allows precise synthesis control over instruction types. With respect to the post-training technique, most methods adopt the SFT Sun et al. (2024b) and RL Huang et al. (2025a); Wang et al. (2024b); Zhang et al. (2024b); Wu et al. (2024b); Ren et al. (2025) for preference alignment. Moreover, training-free workflows have also been proposed to decompose the instructions  Ferraz et al. (2024) and exploit verifiable feedback Wang et al. (2025) for refinement until all the check-boxes are ticked Cook et al. (2024).
@@ 2.3 Chain-of-Thought Reasoning
Recent progress on the chain-of-thought (CoT) Wei et al. (2022); Zhang et al. (2024c); Xu et al. (2025); Besta et al. (2024); Yao et al. (2023); Yang et al. (2024b) has attracted attentions for advancing cognitive capabilities LLMs. Most of these approaches prompted LLMs to break down difficult questions into multiple small units to tackle them systematically before jumping to final answers Kojima et al. (2022); Press et al. (2022); Arora et al. (2022); Zhou et al. (2022); Khot et al. (2022); Zhang et al. (2022). The emergence of OpenAI o1 Jaech et al. (2024) and DeepSeek R1 Guo et al. (2025) has incubated test-time compute scaling techniques for long CoT, where deep reasoning with extensive exploration and feasible reflection are encouraged Chen et al. (2025). Both Monte Carlo Tree Search and RL with process and outcome rewards Qin et al. (2024c); Huang et al. (2024, 2025b) are also emphasized in previous academic attempts to replicate slow-thinking LLMs. Compared with existing studies that focus on the mathematic and logic problem-solving, we are aimed at advancing instruction following capabilities. Motivated by R1 Guo et al. (2025), we incentivize reasoning with group relative policy optimization (GRPO) Shao et al. (2024) and demonstrate that RL with rule-centric rewards ultimately pushes the limits of LLMs in following instructions.
@@ 3.2 Self-Evolving Instructions with Complex Rules and Constraints
We start by selecting a set of seed instructions $D_{\text{seed}}$ from the commonly used WildChat Zhao et al. (2024) and Alpaca Taori et al. (2023) datasets. To ensure the diversity of $D_{\text{seed}}$, we follow Lu et al. (2023) to tag each instruction by its topics and tasks for a wide-range selection of task abilities. Details on the tagging and selection process can be found in Sec. A.3.1.
@@
We adapt self-instruct Wang et al. (2022a) under different fine-grained rules and constraints Zhang et al. (2024a). In view of the verification techniques, both code-execution Qiao et al. (2023) and LLM-as-a-Judge Zheng et al. (2023) are involved to provide evaluation feedback. For the former, we prepare a collection of constraint templates and their executable codes. Then, we randomly instantiate a combination of rules and constraints from the pool. To ensure their mutual compatibility, a pre-defined validity check is enforced to eliminate conflicts (e.g., The first paragraph must start with... and Wrap your entire reply with...). For the latter, we construct pairs of sub-instructions and scoring questions that stress on the style and semantic constraints. Such sub-instructions correspond to constraints with LLM-based evaluation (e.g., Is the answer written in the tone of Confucious?), which are complementary to those relying on code-based evaluation. An off-the-shelf LLM is utilized to perform few-shot in-breadth evolution. With respect to the composition of these atomic sub-instructions, we refer to He et al. (2024b) for definitions of And, Chain, and Selection. These sub-instructions are assembled for the integrated instructions. Details can be found in the Sec. A.3.2.
@@ 3.3 Reinforcement Learning for Reasoning Under Complex Instructions
We propose to incentivize reasoning of LLMs via RL Ouyang et al. (2022); Guo et al. (2025). The development of CoT is optimized towards being structured and sophisticated, which ultimately leads to improved final answers. Without loss of generality we adopt the GRPO Shao et al. (2024) algorithm. Given a query $x$ sampled from the distribution $D$ of complex instructions, the policy model $\pi_{\theta_{\text{old}}}$ from the previous iteration generates a group of $G$ individual outputs $\{y^{i}\}_{i=1}^{G}$. GRPO updates the policy $\pi_{\theta}$ by maximizing the objective:
@@
Compared with the proximal policy optimization (PPO) Schulman et al. (2017), the advantage of the $i$-th output response is computed by normalizing the group-level rewards $\{R^{i}\}_{i=1}^{G}$. PPO optimizes an additional critic model as value function via the generalized advantage estimation (GAE) Schulman et al. (2016). In consideration of the simplicity and our available computing resources, we employ the GRPO as our scale-up RL settings.
@@
To explicitly distinguish the reasoning from the answer contents, we employ a simple minimalist rule-based format reward that checks the existence of ”<think>”, ”</think>”, ”<answer>” and ”</answer>” tags. The format reward encourages the thinking contents encolosed solely between ”<think>” and ”</think>”. It is noted that for R1-series reasoning models, ”<answer>” and ”</answer>” tags are not necessary for answer extraction.
$R^{i}_{\text{format}}=\left\{\begin{array}[]{lcl}+1&&{\textrm{if}\ \mathds{1}(\texttt{<think>}...\texttt{</think>}\texttt{<answer>}...\texttt{</answer>}\in y^{i}),}\\ -1&&{\textrm{otherwise.}}\end{array}\right.$ (4)
@@
$R^{i}_{\text{accuracy}}=\left\{\begin{array}[]{lcl}+2&&{\textrm{if}\ \text{is\_followed}(y^{i}|c_{j})\ \forall c_{j}\in x_{C}^{\text{active}},}\\ \sum_{j=1}^{C}\frac{\mathds{1}(c_{j}\in x_{C}^{\text{active}}\ \&\&\ \text{is\_followed}(y^{i}|c_{j}))}{\mathds{1}(c_{j}\in x_{C}^{\text{active}})}&&{\text{elif}\ \text{is\_followed}(y^{i}|c_{j})\ \exists c_{j}\in x_{C}^{\text{active}},}\\ -2&&{\textrm{otherwise.}}\end{array}\right.$ (5)
It is noted that the detailed implementation of the verification depends on the constraints Zhang et al. (2024a). For the lexical-level constraints (e.g., keywords and phrases), the numerical constraints (e.g., letters and words), and the format constraint (e.g., JSON, XML, LaTeX, HTML, and Markdown), we resort to simple heuristics with python which provides precise feedback Zhou et al. (2023); Dong et al. (2024b). In contrast, for the semantic-level constraints (e.g., themes and perspectives), stylistic constraint (e.g., writing styles, tones, and role-plays), linguistic constraints (e.g., dialects and morphologies), we bring in the reward model $r_{\phi}$ for judgment He et al. (2024b); Qin et al. (2024a); Wen et al. (2024). However, different from the original GRPO where $r_{\phi}$ implicitly scores the responses in terms of instruction following, we explicitly employ $r_{\phi}$ to check the following conditions of constraints. Besides, $r_{\phi}$ delivers scalar scoring in previous studies while we request for efficient boolean validation (True or False). Since multiple constraints might be active for steering generation, we develop the piecewise reward that promotes more constraint satisfaction while penalizing greatly the extreme cases. In total, our rule-centric reward is defined below.
@@
$\mathds{1}(\max(\{R^{i}_{\text{accuracy}}\}_{i=1}^{G})\geq\min(\{\hat{R}^{i}_{\text{accuracy}}\}^{G}_{i=1})\cdot\mathcal{J}_{\text{GRPO}}^{i}.$